# Intermittent time series forecasting with Gaussian Processes and Tweedie likelihood

### v4

@@ 4.5 Ablation study
We compare TweedieGP against a GP model trained with the approximate Tweedie likelihood with $A{\color[rgb]{0.,0.,0.}(y)}=\phi=1$ (implied by the Tweedie loss) and a TweedieGP trained with no scaling. We keep the same experimental setup of the previous section and we show the results in Tab. 4.
@@
$\mathrm{sQ}_{0.8}$ 1.45 1.45 1.44
$\mathrm{sQ}_{0.9}$ 1.26 1.29 1.25
$\mathrm{sQ}_{0.95}$ 1.16 1.23 1.15
$\mathrm{sQ}_{0.99}$ 1.17 1.38 1.15
@@
$\mathrm{sQ}_{0.5}$ 2.35 2.42 2.34
$\mathrm{sQ}_{0.8}$ 2.29 2.31 2.25
$\mathrm{sQ}_{0.9}$ 2.26 2.25 2.20
$\mathrm{sQ}_{0.95}$ 2.35 2.30 2.22
$\mathrm{sQ}_{0.99}$ 3.39 3.42 2.97
@@
OnlineRetail rmsse 0.95 0.95 0.95
$\mathrm{sQ}_{0.5}$ 1.20 1.17 1.19
$\mathrm{sQ}_{0.8}$ 1.34 1.32 1.29
@@
Auto rmsse 0.82 0.82 0.82
$\mathrm{sQ}_{0.5}$ 1.10 1.11 1.11
@@
$\mathrm{sQ}_{0.9}$ 1.16 1.13 1.13
$\mathrm{sQ}_{0.95}$ 1.22 1.19 1.19
$\mathrm{sQ}_{0.99}$ 1.59 1.59 1.55
@@
Carparts rmsse 0.61 0.61 0.61
$\mathrm{sQ}_{0.5}$ 1.04 1.03 1.00
$\mathrm{sQ}_{0.8}$ 1.10 1.20 1.01
$\mathrm{sQ}_{0.9}$ 1.23 1.21 1.14
$\mathrm{sQ}_{0.95}$ 1.36 1.27 1.26
$\mathrm{sQ}_{0.99}$ 2.43 2.28 2.09
@@
Finally, the performance of the approximated Tweedie likelihood with $\phi=A=1$ confirms what discussed in Sec. 3.2.1: the Tweedie loss might produce effective point forecasts by penalizing values far from zero, but the shorter tails negatively affect the estimate of the higher quantiles.
@@ 5 Conclusions
We test two variants of GP with different likelihoods: the negative binomial and the Tweedie. The two models perform similarly, but TweedieGP performs better on the high quantiles. Our model is the first probabilistic model for intermittent time series which adopts the Tweedie likelihood, which is flexible but whose density is difficult to evaluate. Hence we find it useful to publicly release the Tweedie likelihood, making it possible to train novel models probabilistic models for intermittent time series in the future.
Our GP model uses only a RBF kernel; in future works, we could add also a periodic kernel to model the effect of seasonality, which is an open problem in intermittent time series (Fildes et al., 2022). Adding a periodic kernel to the RBF already proved to be successful in smooth time series (Corani et al., 2021). We also leave for future works the comparison between local and global models for intermittent time series.
@@ Appendix A Learning a sparse variational GP
The posterior over latent function $p(\mathbf{f}_{1:T}\mid y_{1:T})$ is not available analytically therefore we need to approximate it. Moreover, in order to optimise the hyperparameters of the model, we also need to approximate the marginal likelihood $p(\mathbf{y}_{1:T})$. We follow Hensman et al. (2015) and use a sparse GP model with inducing points.
@@
The joint distribution of data and latent variables thus becomes $p(\mathbf{y}_{1:T},\mathbf{f}_{1:T},\mathbf{u})=p(\mathbf{y}_{1:T}\mid\mathbf{f}_{1:T})p(\mathbf{f}_{1:T}\mid\mathbf{u})p(\mathbf{u})$. We consider the approximate distribution $q(\mathbf{u})=\mathcal{N}(\mathbf{u}\mid\mathbf{m},\mathbf{S})$, where $\mathbf{m},\mathbf{S}$ are free parameters to be optimized.
[…]
@@
We implemented the log-likelihood function $\log p(\mathbf{y}_{1:T}\mid\mathbf{f})$ for the Tweedie likelihood and then used GPyTorch (Gardner et al., 2018) for optimizing the ELBO.
@@
note that this integral has an analytic solution because the distributions are both Gaussian therefore $\mathbf{f}_{T+1:T+h}$ is a multivariate Gaussian distribution with known mean and covariance; see, e.g., Hensman et al. (2015) for detailed formulas.
@@
In our experiments we choose the number of inducing points as follows. On short time series ($T\leq 200$), we use $T$ inducing points and initialize their locations in correspondence of the training inputs. On longer time series ($T>200$) we use $200$ inducing points, sampling their initial locations from a multinomial distribution with $p(i)\propto\log\left(1+\frac{i}{T}\right)$, $i=1,\ldots,T$. Thus the recent observations, which are more relevant for forecasting, have higher probability of being chosen as initial location of the inducing points. Since the method has a quadratic cost in $m$, the number of inducing points, and it is linear in $T$, the size of the training set, then we achieve quadratic cost in $T$ for $T<200$ and linear cost in $T$ for $T>200$.
@@ B.1 Truncating the infinite summation
$p(y\mid\mu,\phi,\rho)=\begin{cases}\exp(-\frac{\mu^{2-\rho}}{\phi(2-\rho)})&\text{if}\ y=0;\\ A(y)\cdot\exp\left[\frac{1}{\phi}\left(y\frac{\mu^{1-\rho}}{1-\rho}-\frac{\mu^{2-\rho}}{2-\rho}\right)\right]&\text{otherwise}\end{cases}$ (19)
@@
Dunn and Smyth (2005) evaluate $A(y)$ by approximating the infinite sum in eq. (20) with a finite one, retaining its largest elements. We start by finding
@@
then
$V(j)=\frac{z^{j}}{j!\Gamma(\alpha j)}$
and
@@
Stirling’s approximation allows us to simplify the evaluation of the Gamma function: $\Gamma(x+1)\simeq\sqrt{2\pi x}\cdot x^{x}\cdot e^{-x}$. Therefore
@@
$\log V(j)\simeq$ $j\log z-\frac{1}{2}\log 2\pi-\frac{1}{2}\log j-j\log j+j-\frac{1}{2}\log 2\pi$ (24)
@@
$=$ $j\left(\log z+(1+\alpha)-\alpha\log\alpha-(1+\alpha)\log j\right)$ (25)
$-\log 2\pi-\frac{1}{2}\log\alpha-\log j.$ (26)
Differentiating with respect to $j$ we have
@@
$\simeq$ $\log z-\log j-\alpha\log(\alpha j)$ (27)
which is a monotone decreasing function of $j$. For this reason, $\log V$ is a convex function (and therefore $V$ too). To identify its maximum point, we equate it to 00 obtaining
@@
$\log V(j_{\mathrm{max}})=j_{\mathrm{max}}(1+\alpha)-\log 2\pi-\log j_{\mathrm{max}}-\frac{1}{2}\log\alpha$ (30)
From this point, one can start evaluating $V(j)$ for $j=j_{\mathrm{max}}+1,j_{\mathrm{max}}+2,\dots$, until a value $j_{U}$ is found, such that
$\frac{V(j_{\mathrm{max}})}{V(j_{U})}\geq e^{37}$
To compute that, one can either evaluate the difference between eqs. (22) and (30) or compute, again using eq. (22)
@@
where $C_{W}=\log z+(1+\alpha)-\alpha\log\alpha$ does not depend on $j$; such expression can be computed efficiently. Similarly, one can identify $j_{L}$ such that
@@
$r_{L}=\exp\left(\frac{\partial\log V(j)}{\partial j}\right)\biggr{\rvert}_{j=j_{L}-1}\quad\mathrm{and}\quad r_{U}=\exp\left(\frac{\partial\log V(j)}{\partial j}\right)\biggr{\rvert}_{j=j_{U}+1}$
The threshold $e^{-37}$ has been proposed since $e^{-37}\simeq 8\cdot 10^{-17}$ guarantees an appropriate precision using 64-bit floating points. See Dunn and Smyth (2005) for more. In practice, we are interested in computing the log-likelihood
@@
which similarly requires to pass $A(y)$ trough the logarithm. To efficiently implement the evaluation of $A$, it sufficient to refer to eqs. (20), (22), (29), (30), and (31).
@@
Table 5: $j_{U}-j_{L}+1$ for different choices of $y$, $\phi$ and $\rho$. This is the amount of terms used in the approximation of the summation. It is an increasing function with respect to $y$ and $\rho$, decreasing with respect to $\phi$.
Figure 7: The values of $\phi$ and $\rho$ on the fitted TweedieGP on different data sets. Density curves have been estimated via kernel density estimation.
Table 5 shows the amount of terms in the approximation of the summation $A$ is generally affordable; indeed scaling the data by median demand size, the value of $y$ typically lies between 0.5 and 2, while the density plots from Fig. 7 show that often $\phi\in(0.5,5)$ with mode around 1, and $\rho\in(1,1.4)$, with mode around $1.1$.
@@ B.2 Comparison with Tweedie loss
In order to consider the crucial role of $A$ in fitting appropriately a Tweedie likelihood to the data, consider again the Tweedie loss, characterised by $\phi=A=1$: in this simplified version, the implied likelihood in eq. (13) is no longer a probability distribution, as it does not integrate up to $1$. If constrained to do so determining $c:=c(\mu,\phi,\rho)$ such that
[…]
including the normalization constant $c=\frac{\mu^{1-\rho}}{\rho-1}\exp\left(\frac{\mu^{2-\rho}}{2-\rho}\right)$ leads to
@@
Figure 8: In this plot, $\phi=1$. In black, the Tweedie likelihood; the unnormalized density and its normalized version are the dashed lines in red and cyan respectively. The regularizing factor in the unnormalized density makes the tails shorter.
However, the resulting loss is no longer bimodal, as shown in Fig. 8, and empirical results show that having a mass spike in zero comes with the constraint of having short tails. For this reason its performance is not satisfactory on high quantiles, despite the use of the median demand scaling.
@@ Appendix C Alternative results
$\mathrm{sQ}_{0.8}$ 1.74 1.77 1.46 1.65 1.48 1.46
$\mathrm{sQ}_{0.9}$ 1.57 1.65 1.35 1.56 1.29 1.26
$\mathrm{sQ}_{0.95}$ 1.39 1.51 1.37 1.60 1.19 1.16
$\mathrm{sQ}_{0.99}$ 1.25 1.40 1.69 2.25 1.16 1.16
@@
$\mathrm{sQ}_{0.8}$ 2.34 2.33 2.29 2.40 2.27 2.26
$\mathrm{sQ}_{0.9}$ 2.31 2.29 2.23 2.51 2.23 2.20
$\mathrm{sQ}_{0.95}$ 2.33 2.30 2.35 2.75 2.25 2.23
$\mathrm{sQ}_{0.99}$ 3.02 2.98 4.09 4.36 3.22 2.98
@@
OnlineRetail rmsse 0.97 0.98 0.94 0.94 0.94 0.95
$\mathrm{sQ}_{0.5}$ 1.15 1.36 1.19 1.17 1.14 1.14
$\mathrm{sQ}_{0.8}$ 1.30 1.53 1.34 1.39 1.29 1.28
@@
Table 6: Same results as in Tab. 2, but only considering time series with ADI $>1.32$.
We report in Tab. 6 the aggregated metrics considering only time series with ADI $>1.32$. On RAF data set, the scores remain completely unchanged. On M5 and Carparts, the results of all models undergo minimal changes that do not affect the conclusions drawn above.
On OnlineRetail, whose number of time series decreases from over 2400 to 2023, the only noticeable difference is in the performance of WSS on the highest quantile, where it matches that of TweedieGP.
On the Auto data set, whose series with ADI $>1.32$ are 1227, GPs show a slight overall improvement. The performance of WSS and iETS, which model separately demand size and occurrence, generally worsens, especially on the higher quantiles. The results of the empirical quantiles and ADIDAC improve on the lower quantiles, but worsen on the higher ones.

### v5

@@ 4.5 Ablation study
We compare TweedieGP against a GP model trained with the approximate Tweedie likelihood with $A(y)=1$ and $\phi=1$ (implied by the Tweedie loss) and a TweedieGP trained with no scaling. We keep the same experimental setup of the previous section and we show the results in Tab. 4.
@@
$\mathrm{sQ}_{0.8}$ 1.45 1.45 1.45
$\mathrm{sQ}_{0.9}$ 1.26 1.29 1.26
$\mathrm{sQ}_{0.95}$ 1.16 1.23 1.16
$\mathrm{sQ}_{0.99}$ 1.17 1.38 1.16
@@
$\mathrm{sQ}_{0.5}$ 2.36 2.42 2.34
$\mathrm{sQ}_{0.8}$ 2.28 2.30 2.25
$\mathrm{sQ}_{0.9}$ 2.24 2.24 2.19
$\mathrm{sQ}_{0.95}$ 2.32 2.30 2.21
$\mathrm{sQ}_{0.99}$ 3.36 3.40 2.98
@@
OnlineRetail
rmsse 0.95 0.96 0.95
$\mathrm{sQ}_{0.5}$ 1.19 1.17 1.19
$\mathrm{sQ}_{0.8}$ 1.34 1.32 1.30
@@
Auto rmsse 0.81 0.82 0.82
$\mathrm{sQ}_{0.5}$ 1.11 1.11 1.11
@@
$\mathrm{sQ}_{0.9}$ 1.13 1.13 1.13
$\mathrm{sQ}_{0.95}$ 1.18 1.19 1.20
$\mathrm{sQ}_{0.99}$ 1.56 1.59 1.55
@@
Carparts
rmsse 0.61 0.61 0.61
$\mathrm{sQ}_{0.5}$ 1.01 1.03 1.00
$\mathrm{sQ}_{0.8}$ 1.08 1.20 1.02
$\mathrm{sQ}_{0.9}$ 1.15 1.21 1.15
$\mathrm{sQ}_{0.95}$ 1.24 1.27 1.26
$\mathrm{sQ}_{0.99}$ 2.36 2.28 2.09
@@
Finally, the performance of the approximated Tweedie likelihood with $A{\color[rgb]{0.,0.,0.}(y)}=\phi=1$ confirms what discussed in Sec. 3.2.1: the Tweedie loss might produce effective point forecasts by penalizing values far from zero, but the shorter tails negatively affect the estimate of the higher quantiles.
@@ 5 Conclusions
We test two variants of GP with different likelihoods: the negative binomial and the Tweedie. The two models perform similarly, but TweedieGP performs better on the high quantiles. Our model is the first probabilistic model for intermittent time series which adopts the Tweedie likelihood, which is flexible but whose density is difficult to evaluate. Hence we find it useful to publicly release our implementation of the Tweedie likelihood, making it usable to train novel probabilistic models for intermittent time series in the future.
Our GP model uses only a RBF kernel; in future works, we could add also a periodic kernel to model the effect of seasonality, which is an open problem in intermittent time series (FILDES20221283). Adding a periodic kernel to the RBF already proved to be successful in smooth time series (Corani_Benavoli_Zaffalon_2021). We also leave for future works the comparison between local and global models for intermittent time series.
CRediT authorship contribution statement
Stefano Damato: Writing – review & editing, Writing original draft, Visualization, Validation, Software, Resources, Project administration, Methodology, Investigation, Formal analysis, Data curation, Conceptualization. Dario Azzimonti: Writing - review & editing, Writing original draft, Visualization, Validation, Supervision, Resources, Project administration, Methodology, Investigation, Funding acquisition, Formal analysis, Data curation, Conceptualization. Giorgio Corani: Writing - review & editing, Writing original draft, Visualization, Validation, Supervision, Resources, Project administration, Methodology, Investigation, Funding acquisition, Formal analysis, Data curation, Conceptualization.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Acknowledgments
This work is partially funded by the Swiss National Science Foundation (SNF), grant 200021_212164/1. This project has received funding from the European Union’s Horizon Europe Research and Innovation Framework under grant agreement No.101160720.
@@ Appendix A Learning a sparse variational GP
The posterior over latent function $p(\mathbf{f}_{1:T}\mid y_{1:T})$ is not available analytically therefore we need to approximate it. Moreover, in order to optimise the hyper-parameters of the model, we also need to approximate the marginal likelihood $p(\mathbf{y}_{1:T})$. We follow Hensman_Matthews_Ghahramani_2015 and use a sparse GP model with inducing points.
@@
The joint distribution of data and latent variables thus becomes $p(\mathbf{y}_{1:T},\mathbf{f}_{1:T},\mathbf{u})\allowbreak=p(\mathbf{y}_{1:T}\mid\mathbf{f}_{1:T})p(\mathbf{f}_{1:T}\mid\mathbf{u})p(\mathbf{u})$. We consider the approximate distribution $q(\mathbf{u})=\mathcal{N}(\mathbf{u}\mid\mathbf{m},\mathbf{S})$, where $\mathbf{m},\mathbf{S}$ are free parameters to be optimized.
In variational inference we assume that the posterior is approximated by $p(\mathbf{f}_{1:T}\mid\mathbf{y}_{1:T})\approx q(\mathbf{f}_{1:T}\mid\mathbf{y}_{1:T})=\int p(\mathbf{f}_{1:T}\mid\mathbf{u})q(\mathbf{u})d\mathbf{u}$. Since $q(\mathbf{u})$ is Gaussian we can solve the integral analytically. We can then bound the marginal log-likelihood $\log p(\mathbf{y}_{1:T})$ with the standard variational bound (Hensman_Fusi_Lawrence_2013)
@@
We implemented the log-likelihood function $\log p(\mathbf{y}_{1:T}\mid\mathbf{f})$ for the Tweedie likelihood and then used GPyTorch (Gardner_Pleiss_Weinberger_Bindel_Wilson_2018) for optimizing the ELBO.
@@
note that this integral has an analytic solution because the distributions are both Gaussian therefore $\mathbf{f}_{T+1:T+h}$ is a multivariate Gaussian distribution with known mean and covariance; see, e.g., Hensman_Matthews_Ghahramani_2015 for detailed formulae.
@@
In our experiments we choose the number of inducing points as follows. On short time series ($T\leq 200$), we use $T$ inducing points and initialize their locations in correspondence of the training inputs. On longer time series ($T>200$) we use $200$ inducing points, sampling their initial locations from a multinomial distribution with $p(i)\propto\log\left(1+\frac{i}{T}\right)$, $i=1,\ldots,T$. Thus the recent observations, which are more relevant for forecasting, have higher probability of being chosen as initial location of the inducing points. Since the method has a quadratic cost in $m$, the number of inducing points, and it is linear in $T$, the size of the training set, the cost scales quadratically in $T$ for $T\leq 200$ and linearly in $T$ for $T>200$. .
@@ B.1 Truncating the infinite summation
$p(y\mid\mu,\phi,\rho)=\begin{cases}\exp(-\frac{\mu^{2-\rho}}{\phi(2-\rho)})&\text{if}\ y=0;\\ A(y)\cdot\exp\left[\frac{1}{\phi}\left(y\frac{\mu^{1-\rho}}{1-\rho}-\frac{\mu^{2-\rho}}{2-\rho}\right)\right]&\text{otherwise}\end{cases}$
@@
Dunn_Smyth_2005 evaluate $A(y)$ by approximating the infinite sum in eq. (19) with a finite one, retaining its largest elements. We start by finding
@@
then $V(j)=\frac{z^{j}}{j!\Gamma(\alpha j)}$ and
@@
Stirling’s approximation simplifies the evaluation of the Gamma function: $\Gamma(x+1)\simeq\sqrt{2\pi x}\cdot x^{x}\cdot e^{-x}$. Therefore
@@
$\log V(j)\simeq$ $j\log z-\frac{1}{2}\log 2\pi-\frac{1}{2}\log j-j\log j+j-\frac{1}{2}\log 2\pi$
@@
$=$ $j\left(\log z+(1+\alpha)-\alpha\log\alpha-(1+\alpha)\log j\right)$
$-\log 2\pi-\frac{1}{2}\log\alpha-\log j.$
Differentiating with respect to $j$ we obtain
@@
$\simeq$ $\log z-\log j-\alpha\log(\alpha j)$
which is a monotone decreasing function of $j$. For this reason, $\log V$ is a convex function (and therefore $V$ too). To identify its maximum point, we equate it to 0 to get
@@
$\log V(j_{\mathrm{max}})=j_{\mathrm{max}}(1+\alpha)-\log 2\pi-\log j_{\mathrm{max}}-\frac{1}{2}\log\alpha.$ (25)
From here, we evaluate $V(j)$ for $j=j_{\mathrm{max}}+1,j_{\mathrm{max}}+2,\dots$ until a stopping index $j_{U}$ is found, such that
$\frac{V(j_{\mathrm{max}})}{V(j_{U})}\geq e^{37}.$
To compute that, one can evaluate the difference between eqs. (21) and (25) as
@@
where $C_{W}=\log z+(1+\alpha)-\alpha\log\alpha$ does not depend on $j$; such expression can be obtained efficiently. Similarly, one can identify $j_{L}$ such that
@@
$r_{L}=\exp\left(\frac{\partial\log V(j)}{\partial j}\right)\biggr\rvert_{j=j_{L}-1}\quad\mathrm{and}\quad r_{U}=\exp\left(\frac{\partial\log V(j)}{\partial j}\right)\biggr\rvert_{j=j_{U}+1}.$
The threshold $e^{-37}$ has been proposed since $e^{-37}\simeq 8\cdot 10^{-17}$ guarantees an appropriate precision using 64-bit floating points. See Dunn_Smyth_2005 for more. In practice, we are interested in computing the log-likelihood
@@
which similarly requires to pass $A(y)$ trough the logarithm. To implement the evaluation of $A(y)$, it sufficient to refer to eqs. (19), (21), (24), (25), and (26).
@@
Table 5: $j_{U}-j_{L}+1$ for different choices of $y$, $\phi$ and $\rho$. This is the amount of terms used in the approximation of the summation in $A(y)$. It is an increasing function with respect to $y$ and $\rho$, decreasing with respect to $\phi$.
Figure 7: The values of $\phi$ and $\rho$ on the fitted TweedieGP on different data sets. Density curves have been computed via kernel density estimation.
Table 5 shows the amount of terms in the approximation of the summation $A(y)$ is generally affordable; indeed scaling the data by median demand size, the value of $y$ typically lies between 0.5 and 2, while the density plots from Fig. 7 show that often $\phi\in(0.5,5)$ with mode around 1.5, and $\rho\in(1,1.5)$, with mode around $1.1$.
@@ B.2 Comparison with Tweedie loss
In order to node the crucial role of $A(y)$ in fitting appropriately a Tweedie likelihood to the data, consider again the formula of the Tweedie loss, characterised by $A(y)=\phi=1$: in this simplified version, the implied likelihood in eq. (13) is no longer a probability distribution, as it does not integrate up to $1$. Therefore we determine $c:=c(\mu,\phi,\rho)$ such that
$\int_{0}^{+\infty}c\cdot\exp\left(y\frac{\mu^{1-\rho}}{1-\rho}-\frac{\mu^{2-\rho}}{2-\rho}\right)dy=1,$
and we include $c=\frac{\mu^{1-\rho}}{\rho-1}\exp\left(\frac{\mu^{2-\rho}}{2-\rho}\right)$ to obtain
@@
Figure 8: In this plot, $\phi=1$. In black, the Tweedie likelihood; the unnormalized density and its normalized version are the dashed lines in red and cyan respectively.
However, the distribution implied by the Tweedie loss is no longer bimodal, as shown in Fig. 8, and empirical results show that having a mass spike in zero comes with the constraint of having short tails. For this reason its performance is not satisfactory on high quantiles, despite the use of the median demand scaling.
@@ Appendix C Alternative results
$\mathrm{sQ}_{0.8}$ 1.74 1.77 1.46 1.65 1.48 1.47
$\mathrm{sQ}_{0.9}$ 1.57 1.65 1.35 1.56 1.30 1.27
$\mathrm{sQ}_{0.95}$ 1.39 1.51 1.37 1.60 1.19 1.17
$\mathrm{sQ}_{0.99}$ 1.25 1.40 1.69 2.25 1.17 1.17
@@
$\mathrm{sQ}_{0.8}$ 2.34 2.33 2.29 2.40 2.25 2.26
$\mathrm{sQ}_{0.9}$ 2.31 2.29 2.23 2.51 2.19 2.19
$\mathrm{sQ}_{0.95}$ 2.33 2.30 2.35 2.75 2.24 2.22
$\mathrm{sQ}_{0.99}$ 3.02 2.89 4.09 4.36 3.10 2.99
@@
OnlineRetail
rmsse 0.97 0.98 0.94 0.94 0.94 0.95
$\mathrm{sQ}_{0.5}$ 1.15 1.36 1.19 1.17 1.13 1.13
$\mathrm{sQ}_{0.8}$ 1.30 1.53 1.34 1.39 1.28 1.28
@@
Table 6: Same results as in Tab. 2, but only considering time series with ADI >1.32>1.32.
We report in Tab. 6 the aggregated metrics considering only time series with ADI >1.32>1.32. On RAF data set, the scores remain completely unchanged. On M5 and Carparts, the results of all models undergo minimal changes that do not affect the conclusions drawn above. On OnlineRetail, the only noticeable difference is in the performance of WSS on the highest quantile, where it matches that of TweedieGP.
On the Auto data set, whose series with ADI >1.32>1.32 are 1227, GPs show a slight overall improvement. The performance of WSS and iETS, which model separately demand size and occurrence, generally worsens, especially on the higher quantiles. The results of the empirical quantiles and ADIDAC improve on the lower quantiles, but worsen on the higher ones.